How many grams of anhydrous remains when 40.0g of KAl(SO4)2 . 12H2O is heated until all water has been liberated?


1
Expert's answer
2021-05-26T07:07:29-0400

Upon heating, the water liberates according to the equation:

KAl(SO4)2*12H2O --> KAl(SO4)2 + 12H2O


The molar mass of KAl(SO4)2*12H2O is 474.39 g/mol

The molar mass of anhydrous KAl(SO4)2 is 258.21 g/mol


Therefore,

"m(KAl(SO_4)_2)=40.0g(KAl(SO_4)_2\cdot12H_2O)\times\frac{1mol((KAl(SO_4)_2\cdot12H_2O))}{474.39g(KAl(SO_4)_2\cdot12H_2O)}\times\frac{1mol(KAl(SO_4)_2)}{1mol(KAl(SO_4)_2\cdot12H_2O)}\times\frac{258.21g(KAl(SO_4)_2)}{1mol(KAl(SO_4)_2)}=21.8g"


Answer: 21.8 g